OpenStudy (anonymous):

Let f(x) = x2 - 3x - 7. Find f(-3).

OpenStudy (sheraz12345):

Just insert the value in the func ... U asked this type of question b4 as well

OpenStudy (anonymous):

I know..

OpenStudy (sheraz12345):

So do u understand the question

OpenStudy (anonymous):

Yes, but I don't know how to solve it.

OpenStudy (sheraz12345):

Look we have a function here f(x)== x2 - 3x - 7 now we have to find f(-3) which means x=-3 .. Now insert x=-3 in the equation

OpenStudy (anonymous):

I got -5=-2x ;/

OpenStudy (sheraz12345):

What are u doing there ? Show me the steps

OpenStudy (anonymous):

-3=-9x-7

OpenStudy (sheraz12345):

U insert -3 in the func .. Thats it .. What more can u ask for.. U are going in wrong direction.. The answer will not contain x

OpenStudy (anonymous):

I said I didn't know how to solve the equation

OpenStudy (sheraz12345):

This is how u will get the answer \[(-3)^{2}-3(-3)-7\]

OpenStudy (anonymous):

-9+9-7=-7

OpenStudy (sheraz12345):

No 9+9-7 Square of - gives +

OpenStudy (anonymous):

So it's -7

OpenStudy (sheraz12345):

Ahhhh .... Its 11 silly

OpenStudy (anonymous):

This is very simple, all you do is plug in -3 into x \[x^2-3x-7\] so... \[(-3)^2-3(-3)-7\] then just solve.. = 9 +9 -7 =18-7 =11

OpenStudy (sheraz12345):

Thats what i'm telling her

OpenStudy (anonymous):

Oh I see I just tried the problem over again thank you

OpenStudy (sheraz12345):

You need to improve ur basics

OpenStudy (skullpatrol):